# Astrocyte Regulated Neuromorphic Central Pattern Generator Control of Legged Robotic Locomotion

### v2

@@ Abstract
Neuromorphic computing systems, where information is transmitted through action potentials in a bio-plausible fashion, is gaining increasing interest due to its promise of low-power event-driven computing. Application of neuromorphic computing in robotic locomotion research have largely focused on Central Pattern Generators (CPGs) for bionics robotic control algorithms - inspired from neural circuits governing the collaboration of the limb muscles in animal movement. Implementation of artificial CPGs on neuromorphic hardware platforms can potentially enable adaptive and energy-efficient edge robotics applications in resource constrained environments. However, underlying rewiring mechanisms in CPG for gait emergence process is not well understood. This work addresses the missing gap in literature pertaining to CPG plasticity and underscores the critical homeostatic functionality of astrocytes - a cellular component in the brain that is believed to play a major role in multiple brain functions. This paper introduces an astrocyte regulated Spiking Neural Network (SNN)-based CPG for learning locomotion gait through Reward-Modulated STDP for quadruped robots, where the astrocytes help build inhibitory connections among the artificial motor neurons in different limbs. The SNN-based CPG is simulated on a multi-object physics simulation platform resulting in the emergence of a trotting gait while running the robot on flat ground. 23.3×23.3$\times$23.3 × computational power savings is observed in comparison to a state-of-the-art reinforcement learning based robot control algorithm. Such a neuroscience-algorithm co-design approach can potentially enable a quantum leap in the functionality of neuromorphic systems incorporating glial cell functionality.
@@ I Introduction
Current brain-inspired engineered systems have primarily focused on the emulation of bio-plausible computational models of neurons and synapses. Incorporation of other cellular units from the brain is lacking. Over the past few years, a growing body of evidence has demonstrated that glial cells, and in particular astrocytes, play an important role in the maintenance and modulation of neuronal dynamics facilitating brain function [3]. This work is inspired by recent theories suggesting that rich temporal dynamics such as sychrony form the dynamical basis of learning and memory and are expected to play a key role in enabling a dynamical view of intelligence for the design of next-generation brain-inspired engineered learning systems (see report based on 6th US/NIH BRAIN Initiative Investigators Meeting [27]). In particular, our work focuses on spinal central pattern generators (CPGs), which are neural circuits that generate spontaneous rhythmic patterns and serves as a model “brain-like” system to design engineered learning platforms, specifically robotic locomotion controllers.
Multiple parallel works have pointed toward the critical role of astrocytes in spinal CPGs. Recent literature has reported that astrocytes in the spinal cord can influence the activity of the spinal motor network. Specifically, the excitation of spinal cord astrocytes causes a decline in fictive locomotion frequency [2]. Neuroscientists have further illustrated that excited astrocytes suppress the rhythmic locomotion activity by releasing ATP, which is hydrolyzed into adenosine (ADO) immediately. Adenosine works as an antagonist of the excitatory neurotransmitter dopamine (DA) in the motor neurons [1]. Also, researchers have observed that, in larval zebrafish, the futile swimming effort can activate radial astrocytes, which suppresses continuous unsuccessful attempts [29]. Such results are consistent with the claim that astrocytes play a homeostatic role in the neural system [41]. Driven by such insights, we investigate the homeostatic regulation role of astrocytes in neuromorphic Spiking Neural Network (SNN) based CPG circuits for gait learning in quadruped robots. Neuromorphic robotic control has been recently gaining attention due to the promise of enabling low-power edge robotics in resource constrained environments [34, 30, 13]. However, the works remain preliminary with application to simple robotic platforms primarily due to overtly simplistic computational modelling of the biological system. This work forges stronger connections with neuroscience computational models of CPG and astrocyte regulated synaptic plasticity to develop an algorithmic learning framework for gait emergence in legged robotic locomotion. The key distinguishing factors of our work against prior proposals are:
@@
(ii) Astrocyte Regulated CPG Plasticity: While the dynamics of structures mediating CPGs have been extensively studied [11, 4], rewiring or plasticity of CPGs remains relatively underexplored [33, 19], especially in terms of understanding the homeostatic role of astrocytes. We demonstrate that astrocytes, working in parallel with local learning rules like reward-modulated Spike Timing Dependent Plasticity (STDP), suppresses over-activated motor neurons directly or indirectly, by creating inhibitory synapses. This enables the emergence of a gait that provides continuous and smooth locomotion, while allowing us to harness the benefits of local, event-driven learning capability of bio-inspired systems when implemented on neuromorphic chips - paving the pathway for edge processing in robotics [36].
@@ II Related Work and Main Contributions
In recent years, global policy optimization for robotic locomotion through reinforcement learning has been widely investigated [14, 38, 9], where a policy network receives a system state vector including the joint positions and velocities, and outputs the driving signal to the motors. The reinforcement learning alleviates the difficulty of tuning complex parameter sets, however, the policy network-driven locomotion usually requires continuous neural network inference for every control time step, which is computationally expensive. Further, training such systems is even more computationally challenging, thereby limiting its applicability for on-chip learning in edge robotic systems with resource constraints.
Another category of robotic locomotion algorithms involves CPG circuits organizing the collaboration among joints or limbs, where one CPG unit controls a single joint or multiple joints. Some literature has applied classical control methods, such as phase oscillators, for the CPG [22, 24, 45]. From the neuromorphic implementation standpoint, there has been some preliminary work aimed at constructing the CPG by using SNNs [34, 30, 13]. However, existing works on SNN CPG uses over-simplified neuron models and network architectures. For example, the flexor and the extensor motor neuron pools are usually simplified into a network of two neurons, where the internal fine structure is neglected. The lack of inclusion of biological details in the CPG architecture has resulted in limited flexibility of the control system, thereby constraining their applicability to mostly simple robotic platforms like hexapod robots in contrast to the more complex design space of real-world robotic locomotion control of more intensively researched models, such as quadruped robots. In this work, we forge stronger connections with theoretical neuroscience to develop a detailed bio-inspired CPG model and show that astrocyte control is instrumental to ensure optimal and stable gait emergence in robotic quadruped locomotion systems. Local learning mediated gait emergence ensures compatibility of our proposed control system with neuromorphic hardware, thereby leading to the potential of enabling real-time, low-power on-chip learning.
Finally, we also note that while recent work has explored the role of astrocytes in neuromorphic computing, it primarily remains limited to self-repair functionalities [42, 25, 26, 31, 17] and more recently self-attention mechanisms [21], working memory [40], energy minimization in neural networks [8], structure learning [15], among others
@@ III-A Quadruped System Overview
The defining feature of a quadruped system (applicable for both biological systems and robots) is the set of four limbs. In the locomotion of a quadruped system, the contact between the ground and the feet of the limbs provides support to the torso and acceleration to the entire animal/robot. In case of a biological system, the lower motor neurons (located in the ventral horn of the spinal cord) directly control the tension of the joint flexors and extensors in the limbs [28]. The dynamics of the joint movement in the gait during locomotion, specifically the torque, angular position, angular velocity and angular acceleration of the joints are determined by the activity of the lower motor neuron themselves and the communication among the lower motor neurons pools; between the upper motor neuron (located in the brainstem) and the lower motor neuron; and from the Golgi tendon organ to the lower motor neurons [6, 10]. The gait locomotion functionality of a cat is still complete even when the cerebrums are removed, which indicates that the locomotion control system of four-limbed vertebrates is largely a local neural circuit in the brainstem and spinal cord [44]. Deriving insights from such neuroscience studies, we develop a bio-inspired CPG controller that models the collaboration among the limbs of a quadruped system. The framework is evaluated for a quadruped robot in a multi-body interaction physics simulation environment. In this paper, the Unitree A1 quadruped robot [43] system is adopted. The dynamics of the robot is simulated in the MuJoCo platform [39].
Figure 1: SNN-based CPG architecture. a, The reciprocal inhibition of flexor and extensor motor neuron pools, where the flexor and extensor motor neuron pools are connected through V1/V2b inhibitory interneurons. For clarity, only 5 motor neurons for each pool are displayed. The symbols for synapses in the following sub-figures are consistent with that in sub-figure a. b, The thigh muscle inhibits its antagonistic muscle in the calf through an inhibitory interneuron (IIN). c, A sample snapshot of spontaneous rhythmic burst pattern output from the reciprocal inhibition structure, under non-feedback condition. The vertical axis represents the number of motor neurons spiking at each time step. d, The motor neuron pool of the thigh sends synapses (solid lines with diamond ends) to the motor neuron pool of the thigh in all other limbs. It is worth noting here that these synapses could be either excitatory or inhibitory. For clarity, only the synapses from the front-left limb are displayed. FR: front-right, FL: front-left, RR: rear-right, RL: rear-left.
@@ III-B Central Pattern Generator
Critical behavior of the gait locomotion of four-limbed animals are: (i) the alternating activity of the flexor and extensor muscles in the limbs [12, 10] and (ii) the phase difference among the limb movements. The alternating excitation of the two muscles of a single joint allows the joint to swing forward and backward periodically, which further converts the limb movement between the swing and stance phases [5]. In the past decade, research has elucidated the alternating excitation mechanism of the flexor and extensor motor neuron pools - reciprocal inhibition [20]. Past research has divided the entire CPG for gait locomotion into locomotion units that individually control the movement of a single limb [10]. Within the locomotion units, the motor neurons to an individual muscle inhibit or excite the motor neurons to other muscles [10]. The phase difference among the limbs usually corresponds to the specific locomotion speed of the animal [20]. Based on the phase difference, the gaits are usually categorized into walking, trotting and bounding [20]. The following sections will explain (a) how a locomotion unit, controlling a single limb, is composed of reciprocally inhibiting spiking motor neuron groups and control the joint movement, and (b) how the connections among all the locomotion units (i.e., inter-limb connection) determine the phase difference among the limbs.
@@ III-B1 Locomotion Unit
Related experiments illustrate that the flexor and extensor muscle of a single joint always excite in opposite phases, i.e., when the flexor muscle bursts, the extensor muscle does not fire, and vice versa [10]. Also, the duration of the burst of both the flexor and extensor is approximately constant, which indicates that an active muscle will spontaneously stop firing after a specific duration of time [10]. Such a burst pattern raises two critical questions: (1) How does the motor neuron burst terminate spontaneously after a specific duration of time? (2) How does the halt of one muscle’s activity trigger the activity of the antagonistic muscle? We derive qualitative inspiration from neuroscience to answer these questions. The continuous firing of a motor neuron will accumulate $\mathrm{Ca^{2+}}$ in the cytoplasm, which later opens the $\mathrm{K^{+}}$ channels [10]. The decrement in the spike frequency of one muscle’s motor neuron pool leads to the decline of inhibition to the antagonistic motor neuron pool. Based on this insight, in this paper, we assume that each motor neuron is receiving continuous background excitatory stimulation, consequently the flexor and the extensor reciprocally compete with each other, which finally forms periodic complementary burst patterns. The structure of such reciprocal inhibiting SNN for a single joint is shown in Figure 1a. Two types of neurons are contained in this architecture: the motor neuron and the V1/V2b inhibitory interneuron. The motor neuron model is Pacemaker Stochastic Leaky-Integrate and Fire (PSLIF), which includes the cytoplasmic $\mathrm{Ca^{2+}}$ dynamics and the opening and closing of the $\mathrm{K^{+}}$ channel. The dynamics of the membrane potential of the motor neuron $i$ in the motor neuron pool can be formulated as:
@@
Here, $v_{\mathrm{motor},i}$ is the membrane potential of neuron $i$ in the pool. $\tau_{\mathrm{motor}}$ is the motor neuron membrane potential decay time constant. $I_{\mathrm{background}}(t,v)=I_{\mathrm{background},0}+k_{\mathrm{background},v}\cdot v_{\mathrm{torso}}$ is the continuous background stimulation. $I_{\mathrm{background},0}$ is the strength of the background stimulation when the magnitude of the velocity of the torso is zero. $k_{\mathrm{background},v}$ is a positive background stimulation increment coefficient, and $v_{\mathrm{torso}}$ is the magnitude of the velocity of the torso. The velocity-related background stimulation will lead to an increasing frequency of the swing of the limbs when the robot is speeding up. The strength of the background stimulation also includes randomness. In the discrete implementation $\mathbf{Noise}(t)=\mathrm{A}_{\mathrm{random}}\cdot U_{[-1,1]}(t)$ where $U_{[-1,1]}(t)$ is a sampling of a random variable whose distribution is $\mathrm{Uniform[-1,1]}$. The random variable is sampled at every time step of the simulation. $\mathrm{A}_{\mathrm{random}}$ is the amplitude of randomness. $C_{\mathrm{K\_chan}}$ is the potential drop rate when the $\mathrm{K^{+}}$ channel is fully on. $\mathbf{Sigmoid}(x)=\frac{1}{1+e^{-x}}$. $S_{\mathrm{K\_chan}}$ is the sensitivity of the $\mathrm{K^{+}}$ channel to the $\mathrm{Ca^{2+}}$ concentration. $[\mathrm{Ca}^{2+}]$ is the cytoplasmic $\mathrm{Ca^{2+}}$ concentration. $\mathrm{Thres}_{\mathrm{Ca}^{2+}}$ is the $[\mathrm{Ca^{2+}}]$ threshold for opening the $\mathrm{K^{+}}$ channel. $w_{\mathrm{V1/V2b}\xrightarrow{}\mathrm{motor}}$ is a negative synapse strength from the V1/V2b interneuron to the motor neurons, and this value is uniformly the same for all the motor neurons. $w_{\mathrm{motor},j\xrightarrow{}i}$ is the intra-pool excitatory synapse strength from motor neuron $j$ to motor neuron $i$. $t_{\mathrm{V1/V2b\_spike}}$ and $t_{\mathrm{motor\_spike},j}$ are the spike timings of V1/V2b interneuron and motor neuron $j$ respectively. The cytoplasmic $\mathrm{Ca^{2+}}$ dynamics is:
@@
where, $[\mathrm{Ca^{2+}}]$ is the concentration of cytoplasmic $\mathrm{Ca^{2+}}$, $\tau_{\mathrm{Ca^{2+}}}$ is the $\mathrm{Ca^{2+}}$ decay time constant, $r_{\mathrm{Ca^{2+}}}$ is the $[\mathrm{Ca^{2+}}]$ unit increment per action potential, $t_{\mathrm{spike}}$ is the time of firing of the motor neuron. Also, the stochasticity of PSLIF neuron is attributed to the randomness of the condition to trigger an action potential. The probability of triggering an action potential by the PSLIF neuron at any time step is:
@@
For each muscle, the corresponding motor neuron pool contains 20 inter-connected motor neurons. The synapses among the motor neurons within a single pool are generated in a spatial manner: the location of each motor neuron is randomly generated in the 3-dimensional region $[0,1]\times[0,1]\times[0,1]$ from a uniform distribution and the synapses between a pair of motor neurons are generated according to the distance between two motor neurons, as shown below:
@@
While previous work has proposed the possible connection of motor neurons in one locomotion unit [10], the mechanical structure of the limbs of the robot used in this paper is slightly different from a common quadruped animal, i.e., the number of joints responsible for locomotion in a single limb and the length ratio between the thigh and calf are different. Hence, the common quadruped locomotion unit architecture cannot be directly applied to the A1 robot. Therefore, we propose a locomotion unit SNN enabling the feet tip to move in a circular path (specific to the A1 robot). The diagram of the locomotion unit SNN is shown in Figure 1b. There is a second pair of motor neuron pools driving the flexor and extensor in the calf. Additionally, the thigh flexor inhibits the calf extensor, and vice versa. The neuron model of the inhibitory interneuron is SLIF, and they share the same model hyperparameters with the V1/V2b interneurons due to similar functionality. At the same time, the thigh flexor and extensors receive inhibitory sensory input when the thigh is approaching the limit position, which allows the transition between the swing and stance phases. For example, the angular range of the front thigh is $[0.6,1.4]$. The front thigh flexor and the front extensor motor neurons will be inhibited when the thigh enters the regions $[0.6,0.65]$ and $[1.35,1.4]$ correspondingly. Given the diagonal inhibition from the motor neuron pools of the thigh to that of the calf, the phase transition of the calf is completely determined by the thigh motor neuron pools. Therefore the limit position inhibition for calf muscles is not necessary.

### v3

@@ Abstract
Neuromorphic computing systems, where information is transmitted through action potentials in a bio-plausible fashion, is gaining increasing interest due to its promise of low-power event-driven computing. Application of neuromorphic computing in robotic locomotion research have largely focused on Central Pattern Generators (CPGs) for bionics robotic control algorithms - inspired from neural circuits governing the collaboration of the limb muscles in animal movement. Implementation of artificial CPGs on neuromorphic hardware platforms can potentially enable adaptive and energy-efficient edge robotics applications in resource constrained environments. However, underlying rewiring mechanisms in CPG for gait emergence process is not well understood. This work addresses the missing gap in literature pertaining to CPG plasticity and underscores the critical homeostatic functionality of astrocytes - a cellular component in the brain that is believed to play a major role in multiple brain functions. This paper introduces an astrocyte regulated Spiking Neural Network (SNN)-based CPG for learning locomotion gait through Reward-Modulated STDP for quadruped robots, where the astrocytes help build inhibitory connections among the artificial motor neurons in different limbs. The SNN-based CPG is simulated on a multi-object physics simulation platform resulting in the emergence of a trotting gait while running the robot on flat ground. $23.3\times$ computational power savings is observed in comparison to a state-of-the-art reinforcement learning based robot control algorithm. Such a neuroscience-algorithm co-design approach can potentially enable a quantum leap in the functionality of neuromorphic systems incorporating glial cell functionality.
@@ I Introduction
Current brain-inspired engineered systems have primarily focused on the emulation of bio-plausible computational models of neurons and synapses. Incorporation of other cellular units from the brain is lacking. Over the past few years, a growing body of evidence has demonstrated that glial cells, and in particular astrocytes, play an important role in the maintenance and modulation of neuronal dynamics facilitating brain function [1]. This work is inspired by recent theories suggesting that rich temporal dynamics such as synchrony form the dynamical basis of learning and memory and are expected to play a key role in enabling a dynamical view of intelligence for the design of next-generation brain-inspired engineered learning systems (see report based on 6th US/NIH BRAIN Initiative Investigators Meeting [2]). In particular, our work focuses on spinal central pattern generators (CPGs), which are neural circuits that generate spontaneous rhythmic patterns and serve as a model “brain-like” system to design engineered learning platforms, specifically robotic locomotion controllers.
Multiple parallel works have pointed toward the critical role of astrocytes in spinal CPGs. Recent literature has reported that astrocytes in the spinal cord can influence the activity of the spinal motor network. Specifically, the excitation of spinal cord astrocytes causes a decline in fictive locomotion frequency [3]. Neuroscientists have further illustrated that excited astrocytes suppress the rhythmic locomotion activity by releasing ATP, which is metabolized into adenosine (ADO) immediately. Adenosine works as an antagonist of the excitatory neurotransmitter dopamine (DA) in the motor neurons [4]. Also, researchers have observed that, in larval zebrafish, the futile swimming effort can activate radial astrocytes, which suppresses continuous unsuccessful attempts [5]. Further, it is reported that such inhibitory effect of astrocytes on swimming in zebrafish is mediated by a signaling pathway involving ATP - adenosine - $\text{A}_{2}$ receptor [6]. Such results are consistent with the claim that astrocytes play a homeostatic role in the neural system [7]. Driven by such insights, we investigate the homeostatic regulation role of astrocytes in neuromorphic Spiking Neural Network (SNN) based CPG circuits for gait learning in quadruped robots. Neuromorphic robotic control has been recently gaining attention due to the promise of enabling low-power edge robotics in resource-constrained environments [8, 9, 10]. However, the works remain preliminary with application to simple robotic platforms primarily due to overtly simplistic computational modeling of the biological system, such as single-neuron rhythm generators. This work forges stronger connections with neuroscience computational models of CPG and astrocyte regulated synaptic plasticity to develop an algorithmic learning framework for gait emergence in legged robotic locomotion. The key distinguishing factors of our work against prior proposals are:
@@
(ii) Astrocyte Regulated CPG Plasticity: While the dynamics of structures mediating CPGs have been extensively studied [11, 12], rewiring or plasticity of CPGs remains relatively underexplored [13, 14], especially in terms of understanding the homeostatic role of astrocytes. This work illustrates that astrocytes, working in parallel with local learning rules like reward-modulated Spike Timing Dependent Plasticity (STDP), suppress over-activated motor neurons directly or indirectly, by creating inhibitory synapses. This enables the emergence of a gait that provides continuous and smooth locomotion, while allowing us to harness the benefits of local, event-driven learning capability of bio-inspired systems when implemented on neuromorphic chips - paving the pathway for edge processing in robotics [15].
Additionally, the proposed algorithm is neuromorphic hardware compatible. For instance, stochastic leaky-integrate-fire spiking neuron characteristics used herein can be emulated by magnetic tunneling junction spintronic devices [16]. Additionally, the astrocyte models considered in this work are all local models, which contain only local variables and communicate with only the neighboring components in the architecture. As a result, the proposed algorithm is both bio-inspired and compatible with existing neuromorphic hardware or can be implemented by simple extensions of current neuromorphic computing architectures [17].
@@ II Related Work and Main Contributions
In recent years, global policy optimization for robotic locomotion through reinforcement learning has been widely investigated [18, 19, 20], where a policy network receives a system state vector including the joint positions and velocities, and outputs the driving signal to the motors. The reinforcement learning alleviates the difficulty of tuning complex parameter sets. However, the policy network-driven locomotion usually requires continuous neural network inference for every control time step, which is computationally expensive. Further, training such systems is even more computationally challenging, thereby limiting their applicability for on-chip learning in edge robotic systems with resource constraints.
Another category of robotic locomotion algorithms involves CPG circuits organizing the collaboration among joints or limbs, where one CPG unit controls a single joint or multiple joints. Some literature has applied classical control methods, such as phase oscillators, for the CPG [21, 22, 23]. From the neuromorphic implementation standpoint, there has been some preliminary work aimed at constructing the CPG by using SNNs [8, 9, 10]. However, existing works on SNN CPG use over-simplified neuron models and network architectures. For example, the flexor and the extensor motor neuron pools are usually simplified into a network of two neurons, where the internal fine structure is neglected. The lack of inclusion of biological details in the CPG architecture has resulted in limited flexibility of the control system, thereby constraining their applicability to mostly simple robotic platforms like hexapod robot, in contrast to the more complex design space of real-world robotic locomotion control of more intensively researched models, such as quadruped robots. In this work, we forge stronger connections with theoretical neuroscience to develop a detailed bio-inspired CPG model and show that astrocyte control is instrumental to ensure optimal and stable gait emergence in robotic quadruped locomotion systems. Local learning mediated gait emergence ensures compatibility of our proposed control system with neuromorphic hardware, thereby leading to the potential of enabling real-time, low-power on-chip learning.
Finally, we also note that while recent work has explored the role of astrocytes in neuromorphic computing, it primarily remains limited to self-repair functionalities [24, 25, 26, 27, 28] and more recently self-attention mechanisms [29], working memory [30], energy minimization in neural networks [31], structure learning [32], and so on.
@@ III-A Quadruped System Overview
The defining feature of a quadruped system (applicable to both biological systems and robots) is the set of four limbs. In the locomotion of a quadruped system, the contact between the ground and the feet of the limbs provides support to the torso and acceleration to the entire animal/robot. In case of a biological system, the lower motor neurons (located in the ventral horn of the spinal cord) directly control the tension of the joint flexors and extensors in the limbs [33]. The dynamics of the joint movement in the gait during locomotion, specifically the torque, angular position, angular velocity and angular acceleration of the joints are determined by the activity of the lower motor neuron themselves and the communication among the lower motor neurons pools; between the upper motor neuron (located in the brainstem) and the lower motor neuron; and from the Golgi tendon organ to the lower motor neurons [34, 35]. The gait locomotion functionality of a cat is still complete even when the cerebrums are removed, which indicates that the locomotion control system of four-limbed vertebrates is largely a local neural circuit in the brainstem and spinal cord [36]. Deriving insights from such neuroscience studies, we develop a bio-inspired CPG controller that models the collaboration among the limbs of a quadruped system. The framework is evaluated for a quadruped robot in a multi-body interaction physics simulation environment. In this paper, the Unitree A1 quadruped robot [37] system is adopted. The dynamics of the robot is simulated in the MuJoCo platform [38].
Figure 1 shows the proposed CPG architecture for robotic locomotion, which can be summarized below. A locomotion unit for a single joint contains 2 motor neuron pools: the extensor and the flexor. Through internal connections, the extensor and flexor motor neurons can generate alternating bursts. The extensor and flexor activities are converted into the direction and magnitude of the joint actuator torque of the corresponding joint. Through STDP learning in the inter-limb connection modulated by the astrocyte and reward signal, thigh locomotion limbs inhibit or excite each other, through which the robot’s gait emerges.
Figure 1: SNN-based CPG architecture. a, The reciprocal inhibition of flexor and extensor motor neuron pools, where the flexor and extensor motor neuron pools are connected through V1/V2b inhibitory interneurons. For clarity, only 5 motor neurons for each pool are displayed. The symbols for synapses in the following sub-figures are consistent with those in sub-figure a. b, The thigh muscle inhibits its antagonistic muscle in the calf through an inhibitory interneuron (IIN). c, A sample snapshot of spontaneous rhythmic burst pattern output from the reciprocal inhibition structure, under non-feedback condition. The vertical axis represents the number of motor neurons spiking at each time step. d, The motor neuron pool of the thigh sends synapses (solid lines with diamond ends) to the motor neuron pool of the thigh in all other limbs. It is worth noting here that these synapses could be either excitatory or inhibitory. For clarity, only the synapses from the front-left limb are displayed. FR: front-right, FL: front-left, $R$: rear-right, RL: rear-left.
@@ III-B Central Pattern Generator
Critical behaviors of the gait locomotion of four-limbed animals are: (i) the alternating activity of the flexor and extensor muscles in the limbs [39, 35] and (ii) the phase difference among the limb movements. The alternating excitation of the two muscles of a single joint allows the joint to swing forward and backward periodically, which further converts the limb movement between the swing and stance phases [40]. In the past decade, research has elucidated the alternating excitation mechanism of the flexor and extensor motor neuron pools - reciprocal inhibition [41]. Past research has divided the entire CPG for gait locomotion into locomotion units that individually control the movement of a single limb [35]. Within the locomotion units, the motor neurons to an individual muscle inhibit or excite the motor neurons to other muscles [35]. The phase difference among the limbs usually corresponds to the specific locomotion speed of the animal [41]. Based on the phase difference, the gaits are usually categorized into walking, trotting and bounding [41]. The following sections will explain (a) how a locomotion unit, controlling a single limb, is composed of reciprocally inhibiting spiking motor neuron groups and controls the joint movement, and (b) how the connections among all the locomotion units (i.e., inter-limb connection) determine the phase difference among the limbs.
@@ III-B1 Locomotion Unit
Related experiments illustrate that the flexor and extensor muscles of a single joint always excite in opposite phases, i.e., when the flexor muscle bursts, the extensor muscle does not fire, and vice versa [35]. Also, the duration of the burst of both the flexor and extensor is approximately constant, which indicates that an active muscle will spontaneously stop firing after a specific duration of time [35]. Such a burst pattern raises two critical questions: (1) How does the motor neuron burst terminate spontaneously after a specific duration of time? (2) How does the halt of one muscle’s activity trigger the activity of the antagonistic muscle? We derive qualitative inspiration from neuroscience to answer these questions. The continuous firing of a motor neuron will accumulate $\mathrm{Ca^{2+}}$ in the cytoplasm, which later opens the $\mathrm{K^{+}}$ channels [35]. The decrement in the spike frequency of one muscle’s motor neuron pool leads to the decline of inhibition to the antagonistic motor neuron pool. Based on this insight, in this paper, we assume that each motor neuron is receiving continuous background excitatory stimulation, and consequently, the flexor and the extensor reciprocally compete with each other, which finally forms periodic complementary burst patterns. The structure of such a reciprocal inhibiting SNN for a single joint is shown in Figure 1a. Two types of neurons are contained in this architecture: the motor neuron and the V1/V2b inhibitory interneuron. The motor neuron model is Pacemaker Stochastic Leaky-Integrate and Fire (PSLIF), which includes the cytoplasmic $\mathrm{Ca^{2+}}$ dynamics and the opening and closing of the $\mathrm{K^{+}}$ channel. The dynamics of the membrane potential of the motor neuron $i$ in the motor neuron pool can be formulated as:
@@
Here, $v_{\mathrm{motor},i}$ is the membrane potential of neuron $i$ in the pool. $\tau_{\mathrm{motor}}$ is the motor neuron membrane potential decay time constant. $I_{\mathrm{background}}(t,v)=I_{\mathrm{background},0}+k_{\mathrm{background},v}\cdot v_{\mathrm{torso}}$ is the continuous background stimulation. $I_{\mathrm{background},0}$ is the strength of the background stimulation when the magnitude of the velocity of the torso is zero. $k_{\mathrm{background},v}$ is a positive background stimulation increment coefficient, and $v_{\mathrm{torso}}$ is the magnitude of the velocity of the torso. The velocity-related background stimulation will lead to an increasing frequency of the swing of the limbs when the robot is speeding up. The strength of the background stimulation also includes randomness. In the discrete implementation $\mathbf{Noise}(t)=\mathrm{A}_{\mathrm{random}}\cdot U_{[-1,1]}(t)$ where $U_{[-1,1]}(t)$ is a sampling of a random variable whose distribution is $\mathrm{Uniform[-1,1]}$. The random variable is sampled at every time step of the simulation. $\mathrm{A}_{\mathrm{random}}$ is the amplitude of randomness. $C_{\mathrm{K\_chan}}$ is the potential drop rate when the $\mathrm{K^{+}}$ channel is fully on. $\mathbf{Sigmoid}(x)=\frac{1}{1+e^{-x}}$. $S_{\mathrm{K\_chan}}$ is the sensitivity of the $\mathrm{K^{+}}$ channel to the $\mathrm{Ca^{2+}}$ concentration. $\mathrm{Ca}^{2+}]$ is the cytoplasmic $\mathrm{Ca^{2+}}$ concentration. $\mathrm{Thres}_{\mathrm{Ca}^{2+}}$ is the $\mathrm{Ca^{2+}}]$ threshold for opening the $\mathrm{K^{+}}$ channel. $w_{\mathrm{V1/V2b}\xrightarrow{}\mathrm{motor}}$ is a negative synapse strength from the V1/V2b interneuron to the motor neurons, and this value is uniformly the same for all the motor neurons. $w_{\mathrm{motor},j\xrightarrow{}i}$ is the intra-pool excitatory synapse strength from motor neuron $j$ to motor neuron $i$. $t_{\mathrm{V1/V2b\_spike}}$ and $t_{\mathrm{motor\_spike},j}$ are the spike timings of V1/V2b interneuron and motor neuron $j$ respectively. The cytoplasmic $\mathrm{Ca^{2+}}$ dynamics is:
@@
where, $\mathrm{Ca^{2+}}]$ is the concentration of cytoplasmic $\mathrm{Ca^{2+}}$, $\tau_{\mathrm{Ca^{2+}}}$ is the $\mathrm{Ca^{2+}}$ decay time constant, $r_{\mathrm{Ca^{2+}}}$ is the $\mathrm{Ca^{2+}}]$ unit increment per action potential, $t_{\mathrm{spike}}$ is the time of firing of the motor neuron. Also, the stochasticity of PSLIF neuron is attributed to the randomness of the condition to trigger an action potential. The probability of triggering an action potential by the PSLIF neuron at any time step is:
@@
For each muscle, the corresponding motor neuron pool contains 20 inter-connected motor neurons. The synapses among the motor neurons within a single pool are generated in a spatial manner: the location of each motor neuron is randomly generated in the 3-dimensional region $0,1]\times[0,1]\times[0,1]$ from a uniform distribution and the synapses between a pair of motor neurons are generated according to the distance between two motor neurons, as shown below:
@@
While previous work has proposed the possible connection of motor neurons in one locomotion unit [35], the mechanical structure of the limbs of the robot used in this paper is slightly different from a common quadruped animal, i.e., the number of joints responsible for locomotion in a single limb and the length ratio between the thigh and calf are different. Hence, the common quadruped locomotion unit architecture cannot be directly applied to the A1 robot. Therefore, we propose a locomotion unit SNN enabling the feet tips to move in a circular path (specific to the A1 robot). The diagram of the locomotion unit SNN is shown in Figure 1b. There is a second pair of motor neuron pools driving the flexor and extensor in the calf. Additionally, the thigh flexor inhibits the calf extensor, and vice versa. The neuron model of the inhibitory interneuron is SLIF, and they share the same model hyperparameters with the V1/V2b interneurons due to similar functionality. At the same time, the thigh flexor and extensors receive inhibitory sensory input when the thigh is approaching the limit position, which allows the transition between the swing and stance phases. For example, the angular range of the front thigh is $[0.6,1.4]$. The front thigh flexor and the front extensor motor neurons will be inhibited when the thigh enters the regions $[0.6,0.65]$ and $[1.35,1.4]$ correspondingly. Given the diagonal inhibition from the motor neuron pools of the thigh to those of the calf, the phase transition of the calf is completely determined by the thigh motor neuron pools. Therefore the limit position inhibition for calf muscles is not necessary.